OpenStudy (anonymous):

Write the formula for the geometric sequence. 7, 21, 63, 189… A. an=7∙3n-1 B. an=189∙3n-1 C. an=7∙3n D. an=

OpenStudy (anonymous):

An=7*(3^n) because if n=0 you get just 7, if n=1 you get 7*3, aka 21, etc

OpenStudy (anonymous):

need more clarification?

OpenStudy (e.cociuba):

Just multiply by 3: 7, 21, 63, 189, 567, 1701, 5103... The general formula for the geometric sequence is: a(n) = 7*3^(n-1) So it's A.)

OpenStudy (anonymous):

opps yea, that's the one^

OpenStudy (e.cociuba):

Haah lol I do tht all the time 2! lol :p @sschroeder

OpenStudy (anonymous):

i guess it differs depending on what value of N you want to start with, they do the same action but for the one i selected it means N=0 is the start, where as yours would start at n=1... I still think mine should be right Now that i look at it more

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

I seriously feel like it's actually C not A, assuming you would always start with an n=0 value, but it all depends, they do the same thing. does is define what n starts as?

OpenStudy (e.cociuba):

Ohh ya ur rite srry i accidently did a typo.... It's c.) My answer can prove it.. Srry @sschroeder and @ashleyhelget

OpenStudy (anonymous):

no worries, thank you again

OpenStudy (e.cociuba):

Cuz if u distribute the 1 its going to stay a 3.. So therefore it's C.)

OpenStudy (e.cociuba):

Answer: C.)